How can you write -3+y=-3 in slope intercept form
Answer by Cromlix(4381)

Hi there,
I think you have missed out an x
Either
-3x + y = -3
which would convert to:-
y = 3x - 3
OR
-3 + y = -3x
which would convert to:-
y = -3x + 3
The equation has to be in the
y = mx + c form
Hope this helps :-)